a bullet of mass 100 gram moving with 20 m/s strikes a wooden plank and penetrates upto 20 cm calculate the resistance offered by the wooden plank

Grade:11

2 Answers

Paras Verma
7 years ago
Resistance offered-
F=ma ,where m=100g=0.1kg 
accln is to be determined.
v2=u2+as
as bullet is stopped after 20 cm in plank
v=0 ,u=20m/s ,s=20cm=0.2m
On substituting 02=202+2a*0.2
a=-400/0.4
a=-1000 (- sign represents deceleration)
F=0.1*1000 (only magnitude of accln as we are not told to find F vector)
F=100N=Resistance offered by wooden plank
 
 
Paras Verma
7 years ago
I made a mistake v2=u2+2as
though i wonder how did i manage to substitute properly
Anyhow apologies for the mistake